OpenStudy (anonymous):

Evaluate C(197,195).

myininaya (myininaya):

what is C?

OpenStudy (anonymous):

combination

myininaya (myininaya):

197!/[195!(2!) =(197*196*195!)/(2*195!) =197*196/2